Find an equation of the line containing the given pair of points (-2, -4) and (-9, -5) the equation of the line in slope- intercept form is y =
Answer by stanbon(75887)

Find an equation of the line containing the given pair of points (-2, -4) and (-9, -5) the equation of the line in slope- intercept form is y =
-----------------------
slope = (-4--5)/(-2--9) = 1/7
-------
Form: y = mx+b; solve for "b":
-4 = (1/7)(-2)+b
-4 = -2/7 + b
b = -26/7
----
Equation:
y = (1/7)x -(26/7)
===================
Cheers,
Stan H.